# Sound waves from primordial black hole formations

### v1

@@ II.4 Apparent horizon formation
We use the appearance of an apparent horizon as the criterion for black hole formation. An apparent horizon is defined as a marginally trapped surface, which requires that the expansion of one of the two null geodesic congruences orthogonal to this surface vanishes while the other is negative. Let $k^{\pm}$ denote the outgoing ($+$) and ingoing (--) null vectors, given by
$k_{\mu}^{\pm}=\frac{1}{\sqrt{2}}(A,\pm B,0,0),$ (24)
@@ II.5 Numerical implementation
To solve the Misner-Sharp equations numerically, we discretize the radial coordinate $\bar{r}$ using a fourth-order central difference scheme. Before black hole formation, the computational domain extends from the origin to an outer boundary $r_{f}$, chosen to be much larger than the initial perturbation scale $r_{m}$. A staggered grid is used to avoid the coordinate singularity at the origin. Specifically, we define a uniform grid as
@@
where $N$ is the number of grid points and the grid spacing is $\Delta\bar{r}=r_{f}/R_{H}(N-1/2)$. As in the publicly available code SPriBoSH, all dynamical variables are defined at the grid points. To impose boundary conditions, we introduce two ghost cells at the inner and outer boundaries (denoted as $\bar{r}_{-2}$, $\bar{r}_{-1}$, $\bar{r}_{N+1}$, and $\bar{r}_{N+2}$). At the inner boundary, we require the redefined variables to be even functions of $\bar{r}$, i.e., $\bar{X}(\bar{r}_{-2})=\bar{X}(\bar{r}_{2})$ and $\bar{X}(\bar{r}_{-1})=\bar{X}(\bar{r}_{1})$. At the outer boundary, the solution is matched to the FLRW background, so $\bar{X}(r_{N+1})=\bar{X}(r_{N+2})=1$.
@@
At each time step, the formation of an apparent horizon is monitored according to the criteria in Section II.4. Once an apparent horizon is detected, its location is determined via cubic spline interpolation, and an excision procedure is applied to remove the interior region, thereby preventing numerical breakdown due to singularity formation. This strategy, similar to that in Ref. Escrivà (2020), involves reconstructing the grid with the apparent horizon as the new inner boundary and interpolating the dynamical variables onto the revised grid. The evolution then continues until the next excision.
@@ III.1 Formation of compression waves
$K(r)=\mathcal{A}\bar{K}(r)=\mathcal{A}e^{-(r/r_{m})^{2}},$ (35)
where $r_{m}$ is the initial perturbation scale and the peak amplitude $\mathcal{A}$ is related to the perturbation amplitude $\delta_{m}$ by
$\mathcal{A}=\frac{\delta_{m}}{f(\omega)\bar{K}(r_{m})r_{m}^{2}}=\frac{(5+3\omega)\delta_{m}}{3(1+\omega)r_{m}^{2}}.$ (36)
As demonstrated in Ref. Musco (2019), this curvature perturbation results in a compensated energy density perturbation in which the mass excess of the central overdensity is exactly balanced by a surrounding underdensity that asymptotically approaches the background density.
Without loss of generality, we set $r_{m}=10R_{H}$ to satisfy the long-wavelength approximation. Consequently, the time of horizon crossing is $t_{m}=100t_{0}$, and the horizon mass at $t_{m}$ is $M_{H}=100t_{0}$. The outer boundary is chosen as either $r_{f}=150R_{H}$ or $r_{f}=300R_{H}$, depending on the case. We employ $4000$ grid points and use a logarithmic grid with the parameter $C=1+9/N$. By applying a dichotomy method, we determine the critical perturbation amplitude to be $\delta_{c}\approx 0.49774$, in agreement with the results of Ref. Escrivà (2020). This outcome validates our code, and a convergence test is presented in Appendix A.
For subcritical perturbations, previous studies Musco et al. (2005); Polnarev and Musco (2007) have thoroughly investigated the hydrodynamical evolution. They found that the perturbation initially grows and then bounces back to form a compression wave propagating outwards. When $\delta_{m}$ is very close to $\delta_{c}$, a second, more violent bounce is observed. Our simulation results for $\delta_{m}=0.3\ll\delta_{c}$ and $\delta_{m}=0.495\lesssim\delta_{c}$ (see Fig. 1) are consistent with these findings. For $\delta_{m}=0.495$, the first bounce causes the central density to drop below the background level (as indicated by the green and red lines); the resulting rarefaction region is rapidly refilled (purple line), leading to a second bounce (brown line). In contrast, for $\delta_{m}=0.3$, the central density remains above the background value, and only one bounce occurs.
@@
Fig. 3 displays the long-term evolution of the energy density contrast $\delta\rho\equiv\rho-\rho_{b}$ for the two cases. In the near-critical case, after PBH formation, material near the PBH but outside the apparent horizon bounces to form a thin overdense shell, reminiscent of the behavior observed in subcritical perturbations. This overdense shell is separated from the black hole by a rarefaction region, which effectively prevents further accretion of radiation fluid onto the PBH. The time evolution of the PBH mass for the two cases is shown in Fig. 4. It can be seen that once the overdense shell forms (at approximately $t=50t_{m}$), the PBH mass increases very slowly, reaching a final mass of roughly $2.16$ times its initial value. The energy density in the rarefaction region gradually falls below the background value, and then both the overdense shell and the initial underdensity propagate outwards, forming a compression wave. As the compression wave moves outwards, the rarefaction region dissipates, and the energy density between the PBH and the wave eventually returns to the FLRW background value.
@@
While the radiation-dominated era provides a robust framework for PBH formation, scenarios involving phase transitions or thermal history effects Allahverdi et al. (2020)—such as the QCD crossover Schmid et al. (1999); Jedamzik (1997); Widerin and Schmid (1998); Byrnes et al. (2018); Escrivà et al. (2023) or beyond-Standard-Model physics Escrivà and Subils (2023); Escriva et al. (2024)—can significantly soften the equation of state ($\omega<1/3$). In such cases, the PBH formation thresholds decrease substantially Musco and Miller (2013); Sobrinho et al. (2016); Escrivà et al. (2021) due to reduced pressure gradients, and the generation of compression waves is expected to be more challenging. Our simulation results for $\omega=1/5$ and $\omega=1/10$ are presented in Fig. 5. In both cases, the perturbation amplitudes are chosen such that $\delta_{m}-\delta_{c}\approx 0.00226$, consistent with the near-critical case for $\omega=1/3$. It is evident that the compression wave is weaker for $\omega=1/5$ and disappears for $\omega=1/10$, compared to the radiation-dominated case. This observation is consistent with the expectation that a softer equation of state—with its reduced pressure gradient—hinders pressure forces from overcoming gravity and forming compression waves.
@@
In addition to the fact that the equation of state affects the dynamics of PBH formation, it is well-known that the formation threshold Nakama et al. (2014); Musco (2019); Escrivà et al. (2020, 2021), mass Escrivà and Romano (2021), and abundance Germani and Musco (2019) of PBHs depend on the shape of the curvature perturbations. Therefore, it is of interest to study how the formation of compression waves depends on the perturbation profile. To this end, we follow Ref. Musco (2019) and consider a generalized profile for the curvature perturbation:
$K(r)=\left(\frac{r}{\Delta}\right)^{2\lambda}\mathcal{A}\exp\left[-\frac{1}{2}\left(\frac{r}{\Delta}\right)^{2\beta}\right],$ (37)
where $\Delta$ is related to the perturbation scale $r_{m}$ by
$\Delta=\left(\frac{2(\lambda+1)}{\beta}\right)^{-1/2\beta}r_{m}.$ (38)
Here, $\beta$ and $\lambda$ are parameters that control the shape of the perturbation: $\beta$ determines the steepness of the profile, while $\lambda$ controls the location of the peak. Eq. (35) is recovered when $\beta=1$ and $\lambda=0$. We consider three cases: a steeper profile with $\beta=0.5,\lambda=0$; a flatter profile with $\beta=4,\lambda=0$; and an off-centered profile with $\beta=1,\lambda=4$. In all cases, the perturbation amplitudes are chosen to satisfy $\delta_{m}-\delta_{c}\approx 0.00226$ as well. Fig. 6 shows the time evolution of the energy density contrast for these different profiles. The flatter profile leads to a larger pressure gradient and, therefore, a higher formation threshold than the steeper profile. As can be seen in the top panel of the figure, the flatter profile produces a stronger bounce (as indicated by the red lines), reflecting an enhanced pressure gradient, while the steeper profile results in a wider compression wave. For smaller or larger $\beta$, our code is unable to follow the dynamical evolution for extended periods, and we leave the detailed study of the dependence of the compression wave shape on the perturbation profile for future work. For the off-centered profile, simulations indicate that during the early evolution, matter redistributes to fill the central depression, converging toward a centrally peaked profile (see the bottom-left panel of the figure). Consequently, the resulting compression wave is similar to that of a centrally peaked profile (as shown in the bottom-right panel of the figure).
Figure 6: Time evolution of the energy density contrast for different curvature perturbation shapes: steeper profile ($\beta=0.5,\lambda=0$, top-left panel), flatter profile ($\beta=4,\lambda=0$, top-right panel), and off-centered profile ($\beta=1,\lambda=4$, bottom panels). For the off-centered profile, the early evolution is shown in the bottom-left panel, while the late evolution is shown in the bottom-right panel. In all cases, the perturbation amplitudes are chosen such that $\delta_{m}-\delta_{c}\approx 0.00226$.
@@ III.2 Propagation of sound waves
where the approximation holds when the sound wave propagates far from the PBH. In physical coordinates, the peculiar velocity of the sound wave is $U_{s}=au$ (i.e., the velocity with respect to the conformal time in Ref. Guo et al. (2021)), which corresponds to $U-U_{b}$ in the Misner-Sharp formalism. Therefore, the relation between the peculiar velocity and the energy density contrast is
@@
which is consistent with the result in Minkowski spacetime Deng (2021). Our numerical results, presented in Fig. 7, confirm this relation. The figure shows the time evolution of the peculiar velocity for two perturbation amplitudes, with the gray dashed line representing the energy density contrast divided by $4c_{s}$. It is evident that the two quantities agree well in regions far from the PBH.
@@
Eq. (47a) also reveals the propagation characteristics of sound waves in a flat FLRW universe. First, the amplitude of the sound wave (whether at a peak or trough) decays approximately as the reciprocal of the comoving radius. Second, the comoving radius $r_{s}$—defined as the intersection point of the overdense and underdense shells (point $S$ in Fig. 8)—satisfies $\Delta r_{s}=c_{s}R_{H}\Delta a$. In other words, in the comoving frame, the sound wave propagates at a constant speed:
@@
Figure 8: Schematic illustration of the sound wave profile for near-critical perturbations. Point $A$ denotes the trough of the small rarefaction region and represents the beginning of the overdense shell, while point $B$ marks a zero-crossing of the energy density contrast at the end of the underdense shell. Point $S$ is the intersection point of the two shells, and points $P$ and $T$ denote the peak and trough of the sound wave, respectively.
Third, the sound wave as a whole propagates at the speed of sound in the comoving frame while preserving its shape, so the comoving thickness of the shells remains constant during propagation. Similar to the results of the sound shell model for FOPTs, our companion letter Zeng et al. (2025) demonstrates that the thickness of the sound shell also determines the peak frequency of the gravitational wave spectrum in the PBH formation scenario. We will focus on the thickness of the sound shells in the next subsection.
@@ III.3 Thickness of sound shells
In the following, we focus primarily on the near-critical case ($\delta_{m}=0.5$). In this scenario, the sound wave comprises two distinct shells: an overdense shell and an underdense shell, denoted by the subscripts “$+$” and “--”, respectively.
@@
The timing of collisions between sound shells depends on the mean separation between PBHs. If collisions occur before the compression wave reaches the linear regime, the peak position of the gravitational wave spectrum may be affected. To compare these cases, Fig. 9 illustrates the time evolution of the comoving thickness of the shells, measured from the onset of overdense shell formation (defined as the appearance of an inflection point $A$ in the energy density contrast). It can be seen that the comoving thickness of the overdense shell initially increases rapidly and then decreases gradually, while the comoving thickness of the underdense shell steadily increases as it propagates outwards. Nevertheless, both thicknesses vary only slightly, so it is reasonable to assume that the comoving thicknesses of the sound shells remain effectively constant.
@@
Since the mass of a PBH is of the same order as the horizon mass at the re-entry time $t_{m}$, relating the comoving thickness of the sound shells to the Hubble radius at $t_{m}$ allows one to infer the corresponding PBH mass from the gravitational wave spectrum generated by these waves. To this end, we vary the initial length scale of the perturbation to adjust its horizon re-entry time. Fig. 10 shows the comoving thickness of the overdense shell at $t=400t_{m}$ as a function of the comoving Hubble radius at $t_{m}$, with the black dashed line representing a linear fit. The relation is given by
@@
and our numerical results indicate that $n\approx 2.62$. A similar relation holds for the underdense shell, where $n\approx 5.24$. This relation is employed in our companion letter Zeng et al. (2025) to estimate the present-day gravitational wave peak frequency in terms of PBH mass.
@@ III.4 Energy of sound shells
Refs. Carr et al. (2021); García-Bellido et al. (2021) used energy conservation to estimate the energy of the compression waves, given by:
@@
To quantify the energy of the overdense shell, we use the Misner-Sharp mass, which accounts for the total energy of the spacetime, including both gravitational potential and kinetic energy. By subtracting the energy of the FLRW background, we define:
@@
The energy of the overdense shell is then given by $E_{+}=\Delta M(r_{s})-\Delta M(r_{A})$, while the energy of the underdense shell is defined as $E_{-}=\Delta M(r_{B})-\Delta M(r_{s})$, where $r_{s}$, $r_{A}$, and $r_{B}$ denote the comoving radii of points $S$, $A$, and $B$ in Fig. 8, respectively.
Fig. 11 shows the energy of the overdense shell and the PBH mass at the moment of shell formation as functions of the perturbation amplitude. For $\delta_{m}$ near the critical threshold $\delta_{c}$, the PBH mass follows the scaling law Niemeyer and Jedamzik (1999); Hawke and Stewart (2002); Musco et al. (2005):
@@
where $\mathcal{K}_{+}$, $\sigma_{+}$, and $C_{+}$ are constants. Using the data for $\delta_{m}\leq 0.5$, the fit yields $\sigma\approx\sigma_{+}\approx 0.37$, $\mathcal{K}\approx 5.84$, $\mathcal{K}_{+}\approx-4.70$, and $C_{+}\approx 1.52$. In Fig. 11, the green and red dashed lines represent the scaling law fits for the energy of the overdense shell and the PBH mass, respectively. These fits agree well with the numerical results even for $\delta_{m}>0.5$. This result suggests that the overdense shell contains at most $1.52M_{H}$ of energy. In Ref.Germani and Musco (2019), it was demonstrated that although critical collapse leads to a broad PBH mass spectrum, the majority of PBHs form with masses approximately $0.8M_{H}$. From Fig. 11, we observe that the energy of the overdense shell at this PBH mass is about $0.9M_{H}$, indicating that Eq. (51) underestimates this energy.
@@
where $r_{+}$ and $r_{-}$ are the comoving radii of the overdense peak and the underdense trough, respectively. These expressions align with the energy estimates in Ref. Deng et al. (2018), apart from an additional factor of $1/2$. Our numerical results confirm that these relations hold with a relative error of approximately $O(6\%)$. At late times, since $\delta_{\pm}r_{\pm}\sim\mathrm{const}$, $d_{\pm}\sim\mathrm{const}$, $r_{\pm}\sim\sqrt{t}$, $a\sim\sqrt{t}$, and $\rho_{b}\sim 1/t^{2}$, we find $E_{\pm}\sim\mathrm{const}$, which is consistent with our numerical results. However, if photon diffusion is taken into account, energy from the sound waves will be transferred to the background, leading to $\mu$-distortions in the CMB spectrum Deng (2021).

### v3

@@ II.4 Apparent horizon formation
We use the appearance of an apparent horizon as the criterion for black hole formation. An apparent horizon is defined as a marginally trapped surface, which requires that the expansion of one of the two null geodesic congruences orthogonal to this surface vanishes while the other is negative. Let $k^{\pm}$ denote the outgoing (++) and ingoing (-) null vectors, given by
$k_{\mu}^{\pm}=\frac{1}{\sqrt{2}}(-A,\pm B,0,0),$ (24)
@@ II.5 Numerical implementation
To solve the Misner-Sharp equations numerically, we discretize the radial coordinate $\bar{r}$ using a fourth-order central difference scheme. Before black hole formation, the computational domain extends from the origin to an outer boundary $r_{f}$, chosen to be much larger than the initial perturbation scale $r_{i}$. A staggered grid is used to avoid the coordinate singularity at the origin. Specifically, we define a uniform grid as
@@
where $N$ is the number of grid points and the grid spacing is $\Delta\bar{r}=r_{f}/R_{H}(N-1/2)$. As in the publicly available code SPriBoSH Escriva:2019nsa, all dynamical variables are defined at the grid points. To impose boundary conditions, we introduce two ghost cells at the inner and outer boundaries (denoted as $\bar{r}_{-2}$, $\bar{r}_{-1}$, $\bar{r}_{N+1}$, and $\bar{r}_{N+2}$). At the inner boundary, we require the redefined variables to be even functions of $\bar{r}$, i.e., $\bar{X}(\bar{r}_{-2})=\bar{X}(\bar{r}_{2})$ and $\bar{X}(\bar{r}_{-1})=\bar{X}(\bar{r}_{1})$. At the outer boundary, the solution is matched to the FLRW background, so $\bar{X}(r_{N+1})=\bar{X}(r_{N+2})=1$.
@@
At each time step, the formation of an apparent horizon is monitored according to the criteria in Section II.4. Once an apparent horizon is detected, its location is determined via cubic spline interpolation, and an excision procedure is applied to remove the interior region, thereby preventing numerical breakdown due to singularity formation. This strategy, similar to that in Ref. Escriva:2019nsa, involves reconstructing the grid with the apparent horizon as the new inner boundary and interpolating the dynamical variables onto the revised grid. The evolution then continues until the next excision.
@@ III.1 Formation of compression waves
$K(r)=\mathcal{A}\bar{K}(r)=\mathcal{A}e^{-(r/r_{i})^{2}},$ (35)
where $r_{i}$ is the initial perturbation scale and the peak amplitude $\mathcal{A}$ is related to the perturbation amplitude $\delta_{m}$ by
$\mathcal{A}=\frac{\delta_{m}}{f(\omega)\bar{K}(r_{i})r_{i}^{2}}=\frac{(5+3\omega)\delta_{m}}{3(1+\omega)r_{i}^{2}}.$ (36)
As demonstrated in Ref. 1809.02127, this curvature perturbation results in a compensated energy density perturbation in which the mass excess of the central overdensity is exactly balanced by a surrounding underdensity that asymptotically approaches the background density.
Without loss of generality, we set $r_{i}=10R_{H}$ to satisfy the long-wavelength approximation. Consequently, the time of horizon crossing is $t_{m}=100t_{0}$, and the horizon mass at $t_{m}$ is $M_{H}=100t_{0}$. The outer boundary is chosen as either $r_{f}=150R_{H}$ or $r_{f}=300R_{H}$, depending on the case. We employ $4000$ grid points and use a logarithmic grid with the parameter $C=1+9/N$. By applying a dichotomy method, we determine the critical perturbation amplitude to be $\delta_{c}\approx 0.49774\pm 1\times 10^{-5}$, in agreement with the results of Ref. Escriva:2019nsa. This outcome validates our code, and a convergence test is presented in Appendix LABEL:app:convergence.
For subcritical perturbations, previous studies gr-qc/0412063; gr-qc/0605122 have thoroughly investigated the hydrodynamical evolution. They found that the perturbation initially grows and then bounces back to form a compression wave propagating outwards. When $\delta_{m}$ is very close to $\delta_{c}$, a second, more violent bounce is observed. Our simulation results for $\delta_{m}=0.3\ll\delta_{c}$ and $\delta_{m}=0.495\lesssim\delta_{c}$ (see Fig. 1) are consistent with these findings. For $\delta_{m}=0.495$, the first bounce causes the central density to drop below the background level (as indicated by the green and red lines); the resulting rarefaction region is rapidly refilled (purple line), leading to a second bounce (brown line). In contrast, for $\delta_{m}=0.3$, the central density remains above the background value, and only one bounce occurs.
@@
Fig. 3 displays the long-term evolution of the energy density contrast $\delta\rho\equiv\rho-\rho_{b}$ for the two cases. The PBH forms at $t\approx 26t_{m}$ and $t\approx 8.6t_{m}$, respectively. In the near-critical case, after PBH formation, material near the PBH but outside the apparent horizon bounces to form a thin overdense shell, reminiscent of the behavior observed in subcritical perturbations. This overdense shell is separated from the black hole by a rarefaction region, which effectively prevents further accretion of radiation fluid onto the PBH. The time evolution of the PBH mass for the two cases is shown in Fig. 4. It can be seen that once the overdense shell forms (at approximately $t=50t_{m}$), the PBH mass increases very slowly, reaching a final mass of roughly $2.16$ times its initial value. The energy density in the rarefaction region gradually falls below the background value, and then both the overdense shell and the initial underdensity propagate outwards, forming a compression wave. As the compression wave moves outwards, the rarefaction region dissipates, and the energy density between the PBH and the wave eventually returns to the FLRW background value.
@@
While the radiation-dominated era provides a robust framework for PBH formation, scenarios involving phase transitions or thermal history effects 2006.16182—such as the QCD crossover astro-ph/9807257; astro-ph/9605152; astro-ph/9808142; 1801.06138; 2209.06196 or beyond-Standard-Model physics 2211.15674; 2311.17760—can significantly soften the equation of state ($\omega<1/3$). In such cases, the PBH formation thresholds decrease substantially 1201.2379; 1609.01205; 2007.05564 due to reduced pressure gradients, and the generation of compression waves is expected to be more challenging. Our simulation results for $\omega=1/5$ and $\omega=1/10$ are presented in Fig. 5. In both cases, the perturbation amplitudes are chosen such that $\delta_{m}-\delta_{c}\approx 0.00226$, consistent with the near-critical case for $\omega=1/3$. The PBH forms at $t\approx 46t_{m}$ and $t\approx 66.7t_{m}$, respectively. It is evident that the compression wave is weaker for $\omega=1/5$ and disappears for $\omega=1/10$, compared to the radiation-dominated case. This observation is consistent with the expectation that a softer equation of state—with its reduced pressure gradient—hinders pressure forces from overcoming gravity and forming compression waves.
@@ III.2 Propagation of sound waves
where the approximation holds when the sound wave propagates far from the PBH. In physical coordinates, the peculiar velocity of the sound wave is $U_{s}=au$ (i.e., the velocity with respect to the conformal time in Ref. 2007.08537), which corresponds to $U-U_{b}$ in the Misner-Sharp formalism. Therefore, the relation between the peculiar velocity and the energy density contrast is
@@
which is consistent with the result in Minkowski spacetime 2106.09817. Our numerical results, presented in Fig. 6, confirm this relation. The figure shows the time evolution of the peculiar velocity for two perturbation amplitudes, with the gray dashed line representing the energy density contrast divided by $4c_{s}$. It is evident that the two quantities agree well in regions far from the PBH.
@@
Eq. (45a) also reveals the propagation characteristics of sound waves in a flat FLRW universe. First, the amplitude of the sound wave (whether at a peak or trough) decays approximately as the reciprocal of the comoving radius. Second, the comoving radius of sound waves $r_{S}$—defined as the intersection point of the overdense and underdense shells (point $S$ in Fig. 7)—satisfies $\Delta r_{S}=c_{s}R_{H}\Delta a$. In other words, in the comoving frame, the sound wave propagates at a constant speed:
@@
Figure 7: Schematic illustration of the sound wave profile for near-critical perturbations. Point AA denotes the trough of the small rarefaction region and represents the beginning of the overdense shell, while point $B$ marks a zero-crossing of the energy density contrast at the end of the underdense shell. Point $S$ is the intersection point of the two shells, and points $P$ and $T$ denote the peak and trough of the sound wave, respectively.
Third, the sound wave as a whole propagates at the speed of sound in the comoving frame while preserving its shape. For any specific point on the wave (e.g., where the overdense and underdense shells intersect, or at the peak or trough), its comoving radius $r$ satisfies $\Delta r\approx c_{s}R_{H}\Delta a$. Consequently, the shell comoving thickness, defined as the radial distance between two specific points on the wave, remains nearly constant during propagation. Similar to the results of the sound shell model for FOPTs, our companion letter Zeng:2025law demonstrates that the thickness of the sound shell also determines the peak frequency of the GW spectrum in the PBH formation scenario. We will focus on the thickness of the sound shells in the next subsection.
@@ III.3 Thickness of sound shells
In the following, we focus primarily on the near-critical case ($\delta_{m}=0.5$). In this scenario, the sound wave comprises two distinct shells: an overdense shell and an underdense shell, denoted by the subscripts “++” and “-”, respectively.
@@
The timing of collisions between sound shells depends on the mean separation between PBHs. If collisions occur before the compression wave reaches the linear regime, the peak position of the GW spectrum may be affected. To compare these cases, Fig. 8 illustrates the time evolution of the comoving thickness of the shells, measured from the onset of overdense shell formation (defined as the appearance of an inflection point $A$ in the energy density contrast profile). The numerical uncertainties are estimated by performing simulations with five different grid sizes ($N=2000,3000,4000,5000,6000$). It can be seen that the comoving thickness of the overdense shell initially increases rapidly and then decreases gradually, while the comoving thickness of the underdense shell steadily increases as it propagates outwards. Nevertheless, both thicknesses vary only slightly, so it is reasonable to assume that the comoving thicknesses of the sound shells remain effectively constant.
@@
Although the shell comoving thickness is nearly constant over time, we can expect it to be proportional to the initial perturbation length, which determines the horizon re-entry time $t_{m}$ and the corresponding comoving Hubble radius $1/a(t_{m})H(t_{m})$. On the other hand, since the mass of a PBH is of the same order as the horizon mass at the re-entry time $t_{m}$, relating the comoving thickness of the sound shells to the Hubble radius at $t_{m}$ allows one to infer the corresponding PBH mass from the GW spectrum generated by these waves. To this end, we vary the initial length scale of the perturbation to adjust its horizon re-entry time; the timescale of gravitational collapse then varies accordingly. Fig. 9 shows the comoving thickness of the overdense shell at $t=400t_{m}$ (this is a good reference time, after which the shell comving thickness remains nearly constant) as a function of the comoving Hubble radius at $t_{m}$, with the black dashed line representing a linear fit. The relation is given by
@@
and our numerical results indicate that $n\approx 2.62$. A similar relation holds for the underdense shell, where $n\approx 5.24$. This relation is employed in our companion letter Zeng:2025law to estimate the present-day GW peak frequency in terms of PBH mass.
@@ III.4 Energy of sound shells
Refs. 1904.02129; 1904.11482 used energy conservation to estimate the energy of the compression waves, given by:
@@
To quantify the energy of the overdense shell, we use the Misner-Sharp mass, which accounts for the total energy of the spacetime, including both gravitational potential and kinetic energy. Since the Universe is radiation-dominated, the energy of the sound shells is defined as the excess energy compared to the homogeneous background. By subtracting the energy of the FLRW background, we define:
@@
The energy of the overdense shell is then given by $E_{+}=\int_{r_{A}}^{r_{S}}4\pi R^{2}(\rho-\rho_{b})\mathrm{d}R=\Delta M(r_{S})-\Delta M(r_{A})$, while the energy of the underdense shell is defined as $E_{-}=\int_{r_{S}}^{r_{B}}4\pi R^{2}(\rho-\rho_{b})\mathrm{d}R=\Delta M(r_{B})-\Delta M(r_{S})=-\Delta M(r_{S})$, where $r_{S}$, $r_{A}$, and $r_{B}$ denote the comoving radii of points $S$, AA, and $B$ in Fig. 7, respectively.
Fig. 10 shows the energy of the overdense shell and the PBH mass at the moment of shell formation (defined as the time when an inflection point $A$ appears in the energy density contrast profile) as functions of the perturbation amplitude. We choose this time rather than $t=400t_{m}$ because the timescale of gravitational collapse differs for different values of $\delta_{m}$, whereas $t_{m}$ is fixed. The numerical uncertainties are also included in the figure. For $\delta_{m}$ near the critical threshold $\delta_{c}$, the PBH mass follows the scaling law astro-ph/9901292; Hawke:2002rf; gr-qc/0412063:
@@
where $\mathcal{K}_{+}$, $\sigma_{+}$, and $C_{+}$ are constants. Using the data for $\delta_{m}\leq 0.5$, the fit yields $\sigma\approx\sigma_{+}\approx 0.37$, $\mathcal{K}\approx 5.84$, $\mathcal{K}_{+}\approx-4.70$, and $C_{+}\approx 1.52$. In Fig. 10, the green and red dashed lines represent the scaling law fits for the energy of the overdense shell and the PBH mass, respectively. These fits agree well with the numerical results even for $\delta_{m}>0.5$. This result suggests that the overdense shell contains at most $1.52M_{H}$ of energy. In Ref.1805.04087, it was demonstrated that although critical collapse leads to a broad PBH mass spectrum, the majority of PBHs form with masses approximately $0.8M_{H}$. From Fig. 10, we observe that the energy of the overdense shell at this PBH mass is about $0.9M_{H}$, indicating that Eq. (49) underestimates this energy.
@@
where $r_{+}$ and $r_{-}$ are the comoving radii of the overdense peak and the underdense trough, respectively. These expressions align with the energy estimates in Ref. 1804.10059, apart from an additional factor of $1/2$. Our numerical results confirm that these relations hold with a relative error of approximately $O(6\%)$. At late times, since $\delta_{\pm}r_{\pm}\sim\mathrm{const}$, $d_{\pm}\sim\mathrm{const}$, $r_{\pm}\sim\sqrt{t}$, $a\sim\sqrt{t}$, and $\rho_{b}\sim 1/t^{2}$, we find $E_{\pm}\sim\mathrm{const}$, which is consistent with our numerical results. However, if photon diffusion is taken into account, energy from the sound waves will be transferred to the background, leading to $\mu$-distortions in the CMB spectrum 2106.09817.
Since we consider compensated perturbations and enforce outer boundary conditions matching the FLRW solution, the energy of the PBH must be balanced by the energy of the overdense and underdense shells, satisfying the relation
$E_{+}+E_{-}$ $=-\Delta M(r_{A})$ (54)
$=-\int_{0}^{r_{A}}4\pi R^{2}(\rho-\rho_{b})\mathrm{d}R$ (55)
$=-\int_{r_{\mathrm{PBH}}}^{r_{A}}4\pi R^{2}(\rho-\rho_{b})\mathrm{d}R-\Delta M(r_{\mathrm{PBH}})$ (56)
$\approx-M_{\mathrm{PBH}}.$ (57)